Using the Angle-Angle-Side triangle, A=23°, B=48°, a=14cm.
What is the value for angle C, b, and c?
A.) 109; 26.63cm; 33.88cm
B.) 109; 21.48cm; 40.11cm
C.) 19; 26.63cm; 33.88cm
D.) 19; 21.48cm; 40.11cm

Found 2 solutions by Alan3354, stanbon:
Answer by Alan3354(69443)

find the 3rd angle, then use the Law of Sines.

Answer by stanbon(75887)

Using the Angle-Angle-Side triangle, A=23°, B=48°, a=14cm.
What is the value for angle C, b, and c?
----
Use the Law of Sines::
b/sin(48) = 14/sin(23)
b = 26.63
C = 180-(23+48) = 180-71 = 109degrees
-----
Ans:: A
Cheers,
Stan H.
-------------------
A.) 109; 26.63cm; 33.88cm
B.) 109; 21.48cm; 40.11cm
C.) 19; 26.63cm; 33.88cm
D.) 19; 21.48cm; 40.11cm